# POLICEd RL: Learning Closed-Loop Robot Control Policies with Provable Satisfaction of Hard Constraints

### v1

@@ Abstract
In this paper, we seek to learn a robot policy guaranteed to satisfy state constraints. To encourage constraint satisfaction, existing RL algorithms typically rely on Constrained Markov Decision Processes and discourage constraint violations through reward shaping. However, such soft constraints cannot offer verifiable safety guarantees. To address this gap, we propose POLICEd RL, a novel RL algorithm explicitly designed to enforce affine hard constraints in closed-loop with a black-box environment. Our key insight is to force the learned policy to be affine around the unsafe set and use this affine region as a repulsive buffer to prevent trajectories from violating the constraint. We prove that such policies exist and guarantee constraint satisfaction. Our proposed framework is applicable to both systems with continuous and discrete state and action spaces and is agnostic to the choice of the RL training algorithm. Our results demonstrate the capacity of POLICEd RL to enforce hard constraints in robotic tasks while significantly outperforming existing methods.
@@ 1 Introduction
In this paper, we seek to learn a robot control policy guaranteed to satisfy an affine state constraint in a deterministic but black-box environment. The state space region where this constraint is not satisfied corresponds to an unsafe area that must be avoided by the robot. Our key insight is to transform the state space surrounding this unsafe area into a repulsive buffer region as shown on Figure 2. This buffer is made repulsive by learning a policy whose actions push the robot state away from the unsafe area. To enable analytical verification of the repulsive character of our learned policy, we constrain its outputs to be affine over the buffer region. This key property allows to easily guarantee whether an affine constraint is satisfied.
Our proposed framework is agnostic to the choice of the RL training algorithm and is applicable to both systems with continuous and discrete state and action spaces. Additionally, our approach can accommodate black-box environments by using a local measure of their nonlinearity, which can be numerically estimated. To the best of our knowledge, no other work learns a policy in such a way that the closed-loop system is guaranteed to satisfy a hard constraint.
@@
Following this approach, we establish analytical conditions under which our learned policy guarantees constraint satisfaction. A natural follow-up question arising from these conditions is whether they are so stringent that no constraint-satisfying policy exist. To answer this crucial problem, we transform the question of existence of such a constraint-satisfying policy into a tractable linear problem. We will then demonstrate through a number of numerical examples how our method can be implemented. Then, using a 7DOF robotic arm, we compare the performance of our method with a number of representative baselines and demonstrate that our method outperforms all the baselines both in terms of its constraint satisfaction as well as its expected accumulated reward.
@@ 2.2 Constraints in reinforcement learning
A probabilistic constraint comes with the guarantee of satisfaction with some probability threshold [10] and hence ensures a higher level of safety than soft constraints as shown on Figure 3. For instance, [23] derived policies having a high probability of not violating the constraints by more than a small tolerance. Using control barrier functions, [12] guaranteed safe learning with high probability. Since unknown stochastic environments prevent hard constraints enforcement, [45] proposed to learn generative model-based soft barrier functions to encode chance constraints. Similarly, by using a safety index on an unknown environment modeled by Gaussian processes, [55, 25] established probabilistic safety guarantees.
@@
The focus of our work is at the third safety level described in [10] and illustrated on Figure 3. This level corresponds to inviolable hard constraints. Works [13, 37] both learned safe policies thanks to a differentiable safety layer projecting any unsafe action onto the closest safe action. However, since these safety layers only correct actions, a precise model of the robot dynamics model must be known, which is typically unavailable in RL settings. This limitation is also shared by [38, 49], which require some knowledge of the robot dynamics model to compute either backward reachable sets [38] or control barrier functions [49]. To circumvent any knowledge of the robot dynamics model, [53] used an implicit black-box model but is restricted to collision avoidance problems in 2D. To avoid these limitations, the Lagrangian-based approach of [32] and the model-free approach of [50] require to a priori specify the functional form of the safety certificate and as they approach zero violations, the cost signal for their DNN approximator becomes sparse, for these reasons they cannot guarantee the validity of their safety certificate.
@@ 3 Prior Work
In this section, we review prior work [7] upon which we build our approach. The POLICE algorithm [7] utilizes the spline theory of DNNs [6] to modify the training process of a DNN by guaranteeing that the DNN outputs are strictly affine within a user-specified region as shown on Figure 4.
@@ 4 Framework
Our approach focuses on deterministic dynamics with no model mismatch, although it can be readily extended to uncertain dynamics thanks to robust safe control [14]. Additionally, we consider the robot dynamics model $f$ to be an implicit black-box, meaning that we can evaluate $f$ but we do not have access to any explicit knowledge or analytical form of $f$. This is similar to the online RL setting where $f$ is a simulator or where $f$ is the actual robot.
@@ 5 Constrained Reinforcement Learning
In this section, we establish a method to solve Problem 1. To enforce the satisfaction of affine constraint $Cs\leq d$, we construct a repulsive buffer where $C\dot{s}\leq 0$ around the constraint line $Cs=d$ as illustrated on Figure 2. This repulsive buffer will then guarantee that closed-loop trajectories of the robot cannot breach the constraint. We will establish the analytical safety guarantees of our method in this section and then in the next section, we discuss the implementation details of our approach.
@@ 5.1 Guaranteed satisfaction of hard constraints
We choose a deterministic policy modeled by a POLICEd DNN $\mu_{\theta}:\mathcal{S}\rightarrow\mathcal{A}$. We adopt the framework of [7] as discussed in Section 3. We assume that the activation functions of $\mu_{\theta}$ are continuous piecewise affine functions such as (leaky) ReLU, absolute value, or max-pooling. As mentioned in Section 3, this POLICEd DNN architecture allows the user to specify a polytopic region $\mathcal{R}_{\text{user}}$ of the state space $\mathcal{S}$ where the outputs of $\mu_{\theta}$ are affine. We choose $\mathcal{R}_{\text{user}}=\mathcal{B}$ as illustrated on Figure 2.
@@
Having an affine policy on $\mathcal{B}$, we would like to couple it with affine robot dynamics to obtain a simple constraint enforcement process. However, in general, robot dynamics (2) are nonlinear and are unknown to the robot. We will thus use an affine approximation of the robot dynamics inside buffer $\mathcal{B}$ using the following definition.
@@
Let us now prove that the infimum of all approximation measures is $\varepsilon^{*}$, by first showing that $\varepsilon^{*}$ is their largest lower bound. If $\varepsilon$ is an approximation measure, then it has a triplet $(A,B,c)$ satisfying (7). By definition of $\varepsilon^{*}_{ABc}$ and $\varepsilon^{*}$, we then have $\varepsilon\geq\varepsilon^{*}_{ABc}\geq\varepsilon^{*}$. Thus, $\varepsilon^{*}$ is a lower bound to all approximation measures. Let $\alpha$ be a lower bound to all approximation measures. Since all the $\varepsilon^{*}_{ABc}$ are approximation measures, $\alpha\leq\varepsilon^{*}_{ABc}$ for all $A,B,c$, i.e., $\alpha\leq\varepsilon^{*}$ by definition of $\varepsilon^{*}$. Therefore, $\varepsilon^{*}$ is the largest lower bound of all approximation measures.
@@
where the first inequality is a triangular inequality, the second follows from (7) and (2), and the third ineqaulity follows from (9). Using the convexity of polytope $\mathcal{B}$ of vertices $\{v_{1},\ldots,v_{N}\}=\mathcal{V}\big{(}\mathcal{B}\big{)}$, for all $s\in\mathcal{B}$, there exists $\alpha_{1},\ldots,\alpha_{N}\in\mathbb{R}$ such that $\alpha_{k}\geq 0$, $\sum_{k=1}^{N}\alpha_{k}=1$ and $s=\sum_{k=1}^{N}\alpha_{k}v_{k}$. Going beyond vertices, policy (6) yields
@@
where the inequality comes from (5.1) on each vertex $v_{k}$. For any state $s\in\mathcal{B}$, we have
@@ 5.2 Existence conditions
The intuition behind Remark 2 is that the derivative of the constrained states must be actuated to allow immediate corrective actions. A relative degree $1$ is also required by CBFs [4] and by numerous works in constrained RL [25, 32, 33, 45] as we already discussed in Section 2.3.
@@ 6 Implementation
with actions $a(t)\in\mathcal{A}:=[-1,1]^{2}$, states $s(t)\in\mathcal{S}:=[0,1]^{2}$ and time step $\delta t=0.1$. We assume dynamics (20) to be a black-box from the controller’s perspective. The reward signal is proportional to the distance between the state and a target state located at $(0.9,0.9)$, as illustrated on Figure 5.
@@
We consider the line joining states $(0.4,0.7)$ to (1.,0.7)(1.,0.7)( 1 . , 0.7 ) to form an obstacle. We will guarantee that trajectories do not cross this constraint line from below by placing our buffer right under it, as shown on Figure 5. Note that a policy trained to reach the target will not cross the constraint line from above as that would move the state away from the target as seen on Figure 5. Thus, we only need to prevent crossings from below. The constraint is parameterized by $C=[0,1]$ and $d=0.7$ in the notations of (3), i.e., $s_{2}\leq 0.7$ when $s_{1}\geq 0.4$.
@@ 7.1 Inverted pendulum experiment
We begin by testing POLICEd RL on the OpenAI inverted pendulum environment [9] which uses the MuJoCo physics engine [42]. The inverted pendulum environment has a 4-dimensional observation space with cart position $x$, pendulum angle $\theta$, linear cart velocity $\dot{x}$, and angular velocity $\dot{\theta}$ as illustrated in Figure 6. The action space is the force applied to the cart $a\propto\ddot{x}$.
@@
We train policy $\mu_{\theta}$ by iterating over two phases. The first phase is a standard RL training where the initial state is regularly reset either around the origin or inside the affine buffer. Once a reward threshold is met, the constraint training phase starts. The state is iteratively reset to all the vertices of buffer $\mathcal{B}$ and the trajectory is propagated for a single time step to evaluate whether repulsion condition (15) holds. Only the experiences where (15) is not verified are added to the replay memory with negative rewards in order to promote the respect of (15). After several rounds of these updates, the first training phase resumes and the process begins anew until condition (15) holds everywhere on the vertices of $\mathcal{B}$ and the maximal reward is achieved. We summarize this training process in Algorithm 2.
@@
During training we measure the proportion of buffer $\mathcal{B}$ where repulsion condition (15) is satisfied and report these proportions in Figure 8. Since the POLICEd policy achieves maximal reward while completely satisfying repulsion condition (15) at episode 1180 it stops training, whereas Figure 8 shows that the baseline never succeeds in enforcing (15) over the entire buffer. Moreover, the constraint training phase causes a large drop in the baseline rewards as seen on Figure 7. Our POLICEd policy guarantees the satisfaction of constraint $\dot{\theta}\leq 0$ by enforcing repulsion condition (15), i.e., $\ddot{\theta}\leq 0$. Consequently, it guarantees that trajectories starting from $\dot{\theta}(0)<0$ will never reach any state where $\dot{\theta}(t)\geq 0$.
@@ 7.2 Robotic arm
We further showcase our method on a robotic arm with $7$ degrees of freedom. To implement our approach, we rely on the high-fidelity MuJoCo simulator [42]. We developed a custom environment where the robotic arm aims at reaching a 3D target in with its end-effector while avoiding an unsafe region as illustrated on Figure 10. More specifically, we define the observation space of the environment $s\in\mathbb{R}^{10}$ as the arm’s joint angles $s_{joints}\in\mathbb{R}^{7}$ and the X-Y-Z coordinates of the end-effector $s_{end}\in\mathbb{R}^{3}$. At each timestep, the policy provides a change in joint angles $a\in\mathbb{R}^{7}$. The target is kept constant across all episodes. At the start of each episode, the starting state of the arm is uniformly sampled from the joint space.
Figure 10: Robotic arms tasked to bring their end-effectors to the target (green) while avoiding the unsafe region (red). The POLICEd method uses a buffer region (cyan). The arms are shown in two random configurations.
At each timestep, the robotic arm is assigned a reward function $R(s,a)$ which is composed of four terms: the straight line distance $R_{d}(s)$ between the end-effector and the target, a bonus $R_{b}(s,a)$ for reaching the target, a penalty $R_{\text{inf}}(a)$ for generating an infeasible joint state or exiting the state space, and a penalty $R_{\text{unsf}}(s,a)$ for entering the unsafe area.

### v2

@@ Abstract
In this paper, we seek to learn a robot policy guaranteed to satisfy state constraints. To encourage constraint satisfaction, existing RL algorithms typically rely on Constrained Markov Decision Processes and discourage constraint violations through reward shaping. However, such soft constraints cannot offer verifiable safety guarantees. To address this gap, we propose POLICEd RL, a novel RL algorithm explicitly designed to enforce affine hard constraints in closed-loop with a black-box environment. Our key insight is to force the learned policy to be affine around the unsafe set and use this affine region as a repulsive buffer to prevent trajectories from violating the constraint. We prove that such policies exist and guarantee constraint satisfaction. Our proposed framework is applicable to both continuous and discrete state-action spaces and is agnostic to the choice of the RL training algorithm. Our results demonstrate the capacity of POLICEd RL to enforce hard constraints in robotic tasks while significantly outperforming existing methods. Code available at https://iconlab.negarmehr.com/POLICEd-RL/
@@ 1 Introduction
In this paper, we seek to learn a robot control policy guaranteed to satisfy an affine state constraint in a deterministic but black-box environment. The state space region where this constraint is not satisfied corresponds to an unsafe area that must be avoided by the robot. Our key insight is to transform the state space surrounding this unsafe area into a repulsive buffer region as shown in Figure 2. This buffer is made repulsive by learning a policy whose actions push the robot state away from the unsafe area. To enable analytical verification of the repulsive character of our learned policy, we constrain its outputs to be affine over the buffer region. This key property allows to easily guarantee whether an affine constraint is satisfied.
Our proposed framework is agnostic to the choice of the RL training algorithm but relies on its convergence to guarantee safety. Our method is applicable to both systems with continuous and discrete state and action spaces. Additionally, our approach can accommodate black-box environments by using a local measure of their nonlinearity, which can be numerically estimated. To the best of our knowledge, no other work learns a policy in such a way that the closed-loop system is guaranteed to satisfy a hard constraint after training.
@@
Following this approach, we establish analytical conditions under which our learned policy guarantees constraint satisfaction after training. A natural follow-up question arising from these conditions is whether they are so stringent that no constraint-satisfying policy exist. To answer this crucial problem, we transform the question of existence of such a constraint-satisfying policy into a tractable linear problem. We will then demonstrate through a number of numerical examples how our method can be implemented. Then, using a 7DOF robotic arm, we compare the performance of our method with a number of representative baselines and demonstrate that our method outperforms all the baselines both in terms of its constraint satisfaction as well as its expected accumulated reward.
@@ 2.2 Constraints in reinforcement learning
A probabilistic constraint comes with the guarantee of satisfaction with some probability threshold [12] and hence ensures a higher level of safety than soft constraints as shown in Figure 3. For instance, [30] derived policies having a high probability of not violating the constraints by more than a small tolerance. Using control barrier functions, [14] guaranteed safe learning with high probability. Since unknown stochastic environments prevent hard constraints enforcement, [59] proposed to learn generative model-based soft barrier functions to encode chance constraints. Similarly, by using a safety index on an unknown environment modeled by Gaussian processes, [69, 33] established probabilistic safety guarantees.
@@
The focus of our work is at the third safety level described in [12] and illustrated in Figure 3. This level corresponds to inviolable hard constraints. Works [16, 47] both learned safe policies thanks to a differentiable safety layer projecting any unsafe action onto the closest safe action. However, since these safety layers only correct actions, a precise model of the robot dynamics model must be known, which is typically unavailable in RL settings. This limitation is also shared by [49, 63], which require some knowledge of the robot dynamics model to compute either backward reachable sets [49] or control barrier functions [63]. To circumvent any knowledge of the robot dynamics model, [67] used an implicit black-box model but is restricted to collision avoidance problems in 2D. To avoid these limitations, the Lagrangian-based approach of [41] and the model-free approach of [64] require to a priori specify the functional form of the safety certificate and as they approach zero violations, the cost signal for their DNN approximator becomes sparse, for these reasons they cannot guarantee the validity of their safety certificate.
@@ 2.3 Relative degree of constraints
2.4 Black-box safety with control theory
While CBFs are prevalent in safe RL [42, 48, 59, 64], other approaches backed by control theory have also investigated safety in black-box RL. Most notably, model predictive control (MPC) is particularly well-suited to bring safety guarantees to RL settings [29]. MPC schemes predict safe and optimal control inputs in a receding-horizon fashion by using a model of the dynamics to anticipate their behavior [29]. Implementing MPC for black-box systems typically requires learning either a robust [6, 18] or stochastic [39] model of the dynamics. Then, the theory of robust [9] and stochastic control [56] can provide safety guarantees despite the uncertainty or disturbances in the learned models [29]. Similarly to CBFs, MPC can also serve as safety filter to enforce constraints on learned policies [29, 57]. While robust control tends to be overly conservative in practice [58] and stochastic MPC offers limited safety guarantees [43], the main limitation of black-box MPC is the unavoidable and notoriously computationally difficult receding-horizon optimization problem to be solved at each time step [29].
More closely related to this work is the literature on optimal control of constrained piecewise affine systems [15]. Indeed, nonlinear dynamics can be locally approximated as piecewise affine to develop robust controllers with provable performance guarantees [24, 28] while enforcing linear temporal logic specifications [31]. However, these works assume knowledge of the robot dynamics.
@@ 3 Prior Work
In this section, we review prior work [8] upon which we build our approach. The POLICE algorithm [8] utilizes the spline theory of DNNs [7] to modify the training process of a DNN by guaranteeing that the DNN outputs are strictly affine within a user-specified region as shown in Figure 4.
@@ 4 Framework
Our approach focuses on deterministic dynamics with no model mismatch, although it can be readily extended to uncertain dynamics thanks to robust safe control [17]. Additionally, we consider the robot dynamics model $f$ to be an implicit black-box, meaning that we can evaluate $f$ but we do not have access to the equations or analytical form of $f$. This is similar to the online RL setting where $f$ is a simulator or where $f$ is the actual robot.
@@ 5 Constrained Reinforcement Learning
In this section, we establish a method to solve Problem 1. To enforce the satisfaction of affine constraint $Cs\leq d$, we construct a repulsive buffer where $C\dot{s}\leq 0$ around the constraint line $Cs=d$ as illustrated in Figure 2. This repulsive buffer will then guarantee that closed-loop trajectories of the robot cannot breach the constraint. We will establish the analytical safety guarantees of our method in this section and then in the next section, we discuss the implementation details of our approach.
@@ 5.1 Guaranteed satisfaction of hard constraints
We choose a deterministic policy modeled by a POLICEd DNN $\mu_{\theta}:\mathcal{S}\rightarrow\mathcal{A}$. We adopt the framework of [8] as discussed in Section 3. We assume that the activation functions of $\mu_{\theta}$ are continuous piecewise affine functions such as (leaky) ReLU, absolute value, or max-pooling. As mentioned in Section 3, this POLICEd DNN architecture allows the user to specify a polytopic region $\mathcal{R}_{\text{user}}$ of the state space $\mathcal{S}$ where the outputs of $\mu_{\theta}$ are affine. We choose $\mathcal{R}_{\text{user}}=\mathcal{B}$ as illustrated in Figure 2.
@@
Having an affine policy on $\mathcal{B}$, we would like to couple it with affine robot dynamics to obtain a simple constraint enforcement process. However, in general, robot dynamics (2) are nonlinear. We will thus use an affine approximation of the robot dynamics inside buffer $\mathcal{B}$ using the following definition.
@@
Let us now prove that the infimum of all approximation measures is $\varepsilon^{*}$, by first showing that $\varepsilon^{*}$ is their largest lower bound. If $\varepsilon$ is an approximation measure, then it has a triplet $(A,B,c)$ satisfying (7). By definition of $\varepsilon^{*}_{ABc}$ and $\varepsilon^{*}$, we then have $\varepsilon\geq\varepsilon^{*}_{ABc}\geq\varepsilon^{*}$. Thus, $\varepsilon^{*}$ is a lower bound to all approximation measures.
Let $\alpha$ be a lower bound to all approximation measures. Since all the $\varepsilon^{*}_{ABc}$ are approximation measures, $\alpha\leq\varepsilon^{*}_{ABc}$ for all $A,B,c$, i.e., $\alpha\leq\varepsilon^{*}$ by definition of $\varepsilon^{*}$. Therefore, $\varepsilon^{*}$ is the largest lower bound of all approximation measures.
@@
where the first inequality is a triangular inequality, the second follows from (7) and (2), and the third ineqaulity follows from (9). Using the fact that an affine function is uniquely determined by its values on the vertices of a full-dimensional polytope [26], [28], we will show that (5.1) is also valid all over polytope $\mathcal{B}$ and not just at its vertices $\mathcal{V}\big{(}\mathcal{B}\big{)}$. More specifically, since $\mathcal{B}$ is the convex hull of its vertices $\{v_{1},\ldots,v_{N}\}=\mathcal{V}\big{(}\mathcal{B}\big{)}$ [26], for all $s\in\mathcal{B}$, there exists $\alpha_{1},\ldots,\alpha_{N}\in\mathbb{R}$ such that $\alpha_{k}\geq 0$, $\sum_{k=1}^{N}\alpha_{k}=1$ and $s=\sum_{k=1}^{N}\alpha_{k}v_{k}$. Then, (6) yields
@@
where the inequality comes from (5.1) applied on each vertex $v_{k}$. For any state $s\in\mathcal{B}$, we have
@@ 5.2 Existence conditions
The intuition behind Remark 2 is that the derivative of the constrained states must be actuated to allow immediate corrective actions. A relative degree $1$ is also required by CBFs [4] and by numerous works in constrained RL [33, 41, 42, 59] as we already discussed in Section 2.3. Let us now detail the implementation of our method.
@@ 6 Implementation
with actions $a(t)\in\mathcal{A}:=[-1,1]^{2}$, states $s(t)\in\mathcal{S}:=[0,1]^{2}$ and time step $\delta t=0.1$. We assume dynamics (20) to be a black-box from the controller’s perspective. The reward signal is proportional to the distance between the state and a target state located at $(0.9,0.9)$, as illustrated in Figure 5.
@@
We consider the line joining states $(0.4,0.7)$ to (1.,0.7)(1.,0.7)( 1 . , 0.7 ) to form an obstacle. We will guarantee that trajectories do not cross this constraint line from below by placing our buffer right under it, as shown in Figure 5. Note that a policy trained to reach the target will not cross the constraint line from above as that would move the state away from the target as seen in Figure 5. Thus, we only need to prevent crossings from below. The constraint is parameterized by $C=[0,1]$ and $d=0.7$ in the notations of (3), i.e., $s_{2}\leq 0.7$ when $s_{1}\geq 0.4$.
@@ 7.1 Inverted pendulum experiment
We begin by testing POLICEd RL on the OpenAI inverted pendulum environment [11] which uses the MuJoCo physics engine [53]. The inverted pendulum environment has a 4-dimensional observation space with cart position $x$, pendulum angle $\theta$, linear cart velocity $\dot{x}$, and angular velocity $\dot{\theta}$ as illustrated in Figure 6. The action space is the force applied to the cart $a$ proportional to $\ddot{x}$.
@@
Note that training to satisfy a constraint is often adversarial to accomplishing the task, and thus often causes training instability. Throughout our development process, we discovered techniques to greatly improve training time and sample complexity, which are further discussed in Appendix A. For example, with the inverted pendulum scenario, we train policy $\mu_{\theta}$ by iterating over two phases. The first phase is a standard RL training where the initial state is regularly reset either around the origin or inside the affine buffer. Once a reward threshold is met, the constraint training phase starts. The state is iteratively reset to all the vertices of buffer $\mathcal{B}$ and the trajectory is propagated for a single time step to evaluate whether repulsion condition (15) holds. Only the experiences where (15) is not verified are added to the replay memory with negative rewards in order to promote the respect of (15). After several rounds of these updates, the first training phase resumes and the process begins anew until condition (15) holds everywhere on the vertices of $\mathcal{B}$ and the maximal reward is achieved. We summarize this training process in Algorithm 2.
@@
During training we measure the proportion of buffer $\mathcal{B}$ where repulsion condition (15) is satisfied and report these proportions in Figure 8. Since the POLICEd policy achieves maximal reward while completely satisfying repulsion condition (15) at episode 1180 it stops training, whereas Figure 8 shows that the baseline never succeeds in enforcing (15) over the entire buffer. Moreover, the constraint training phase causes a large drop in the baseline rewards as seen in Figure 7. Our POLICEd policy guarantees the satisfaction of constraint $\dot{\theta}\leq 0$ by enforcing repulsion condition (15), i.e., $\ddot{\theta}\leq 0$. Consequently, it guarantees that trajectories starting from $\dot{\theta}(0)<0$ will never reach any state where $\dot{\theta}(t)\geq 0$.
@@ 7.2 Robotic arm
We further showcase our method on a robotic arm with $7$ degrees of freedom. To implement our approach, we rely on the high-fidelity MuJoCo simulator [53]. We developed a custom environment where the robotic arm aims to reach a 3D target with its end-effector while avoiding an unsafe region as illustrated in Figure 10. More specifically, we define the state space of the environment $s\in\mathbb{R}^{10}$ as the arm’s joint angles $s_{joints}\in\mathbb{R}^{7}$ and the X-Y-Z coordinates of the end-effector $s_{end}\in\mathbb{R}^{3}$. At each timestep, the policy provides a change in joint angles $a\in\mathbb{R}^{7}$. The target is kept constant across all episodes. At the start of each episode, the starting state of the arm is uniformly sampled from the joint space.
Figure 10: Robotic arms tasked to bring their end-effectors to the target (green) while avoiding the constraint surface (red). The POLICEd method uses a buffer region (cyan). The arms are shown in two random configurations.
At each timestep, the robotic arm is assigned a reward function $R(s,a)$ which is composed of four terms: the straight line distance $R_{d}(s)$ between the end-effector and the target, a bonus $R_{b}(s,a)$ for reaching the target, a penalty $R_{\text{inf}}(a)$ for generating an infeasible joint state or exiting the state space, and a penalty $R_{\text{unsf}}(s,a)$ for violating the constraint.